OpenStudy (anonymous):

how do u do: 6 (x - 5)= 4x - 50 ? and what r the steps ?

OpenStudy (jiteshmeghwal9):

6(x-5)=6x-30

OpenStudy (jiteshmeghwal9):

6x-30=4x-50

OpenStudy (jiteshmeghwal9):

now take variables @ one side & constants @ one side as 6x-4x=-50+30 2x=-20 x=-20/2=-10

OpenStudy (anonymous):

thank u :D

OpenStudy (jiteshmeghwal9):

yw :)